# Schrödinger cats escape confinement

### 08 Jan 2024 v1

@@ Abstract
We consider local quenches from initial states generated by a single spin flip in either the true or the false vacuum state of the confining quantum Ising spin chain in the ferromagnetic regime. Contrary to global quenches, where the light-cone behaviour is strongly suppressed, we find a significant light-cone signal propagating with a nonzero velocity besides the expected localised oscillating component. Combining an analytic representation of the initial state with a numerical description of the relevant excitations using the two-fermion approximation, we can construct the spectrum of post-quench excitations and their overlaps with the initial state, identifying the underlying mechanism. For confining quenches built upon the true vacuum, the propagating signal consists of Schrödinger cats of left and right-moving mesons escaping confinement. In contrast, for anti-confining quenches built upon the false vacuum, it is composed of Schrödinger cats of left and right-moving bubbles which escape Wannier-Stark localisation.
@@ 1 Introduction
It was shown in [3] that the non-equilibrium dynamics of the quantum Ising spin chain is radically altered by confinement. In this context, non-equilibrium time evolution was initiated by a quantum quench, i.e., a sudden change in the Hamiltonian [4, 5]. This protocol is of paradigmatic relevance, and today, it is routinely engineered in experiments on closed quantum systems [6, 7, 8, 9, 10, 11, 12, 13]. In a large class of such systems, including various spin chains with short-range interactions or interacting bosonic or fermionic systems, quasi-particles have an upper bound on their velocity [14], leading typically to a distinctive light-cone pattern in time-dependent correlation functions [15, 4, 5, 16, 17, 18, 19, 20]. However, as shown in [3], the confining forces can suppress the light-cone spreading of correlations. Dynamical confinement and its effects has recently been confirmed in numerous systems  [21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@ 2.1 Confinement in the Ising spin chain
$\chi=2Jh_{l}m\,,$ (2.4)
between kinks, which confines them into meson excitations. Choosing for definiteness $m>0$, the true ground state of the system is a perturbation of the one with positive magnetisation $\ket{+}$, while $\ket{-}$ becomes a metastable state, a.k.a. the false vacuum. In the following, we stay within the scenario of strong confinement, which corresponds to a transverse field $h_{t}$ far below its critical value $1$ and consider values of the longitudinal field $h_{l}$ of order $0.1$.
@@
where $\hat{k}_{1,2}$ and $\hat{x}_{1,2}$ are the canonical momentum and coordinate operators for the kinks. The spectrum of this Hamiltonian can be computed either by a numerical solution of the Schrödinger equation or using a simple semi-classical approximation [47]. We avoid using values that are too large for $h_{l}$ to ensure that even the most spatially localised meson excitations have support over several lattice sites, which makes it possible to neglect the short-range deviations from the linear potential. These were computed in [47] and result from the kinks not being exactly point-like: the interpolation between the two vacua takes place over a crossover region of finite length.
@@
We note that when the polarisation of the initial state is opposite to the longitudinal field present after the quench, the time evolution is again localised in the global case; however, in this case, the mechanism behind this effect Wannier-Stark localisation due to Bloch oscillations [28]. This corresponds to starting the dynamics from a false vacuum state, whose decay is prevented by Bloch oscillations [34]. Contrary to confinement, which also occurs in the continuum limit [52], Wannier-Stark localisation is a lattice effect resulting from the dispersion relation’s periodicity over the Brillouin zone.
@@ 2.2 Phenomenology of local quenches
The first quench protocol is the confining quench starting from the initial state
@@
As a first step, the system’s ground state was found by applying the TEBD method with imaginary time evolution and validated by comparing it to the exact diagonalisation results. We then applied the spin flip and simulated the real-time evolution of the system. For the imaginary time evolution, we chose the time step $dt=0.05$, the total number of steps $N=4000$ and the maximum bond dimension $\chi_{\text{max}}=50$. For the real-time evolution, the parameters were set to $dt=0.01$ and $\chi_{\text{max}}=300$. The Schmidt values were truncated at $10^{-8}$ for both imaginary and real-time evolutions. We verified that all results were robust under reasonable variations of these parameters.
@@
Note that while the localised component qualitatively matches the predicted envelope, there is a component forming a prominent light cone which lies outside the red curves. In contrast to the results for global quenches [3], the escaping fronts in local quenches are of the same order of magnitude as the localised component. Our primary interest is to identify the mechanism underlying their presence.
@@
where |$\ket{+}_{h_{t},h_{l}=0}$ is the positively polarised ground state of the pure transverse Hamiltonian. In this case, we take the post-quench longitudinal field $h_{l}$ to be negative, and so the initial state corresponds to a metastable (false vacuum) state for the Hamiltonian (2.1). Note that such quenches are not purely local, i.e. they necessarily involve a global component; however, it turns out that this component is so small that it can be neglected in our subsequent considerations.
@@ 3.1 Confinement: mesons
where $\omega(\hat{k};K)=\varepsilon(K/2+\hat{k})+\varepsilon(K/2-\hat{k})$. From now on, we omit the hats from the operators. Considering an infinite chain with position $x$ taking values in $\mathbb{Z}$, the Schrödinger equation can be written as
@@
The $n$th level of the above Schrödinger problem corresponds to a meson one-particle state $|M_{n}(K)\rangle$ with momentum $K$ and species number $n$, with $E_{n}(K)$ giving the energy of these states relative to the ground state in the two-fermion approximation, which we order to increase with $n$. For any value of $K$, the meson spectrum is bounded from below and unbounded from above.
@@
$d_{\text{WS}}=\frac{\epsilon(\pi)-\epsilon(0)}{|\chi|}=\frac{2h_{t}}{|\chi|}\,,$ (3.8)
@@
up to a phase factor. For the collisionless case, imposing periodicity of the wave function in the Brillouin zone as $\widetilde{\psi}(\pi)=\widetilde{\psi}(-\pi)$ leads to a Bohr-Sommerfeld quantisation condition of the form
@@
This can be checked to match numerically the wave function obtained by explicit solution of (3.2) for large $n$. For the case $h_{t}=0.25$ and $h_{l}=0.1$, the agreement is already excellent for $n=6$ and rapidly improves for larger $n$.
@@ 3.2 Anti-confinement: bubbles
which grows monotonically with $h_{t}$, and is bounded from below by $2/|h_{l}|$. For the transverse and longitudinal fields considered here, the critical bubble size is always more than an order of magnitude larger than the lattice spacing, which plays an important role in suppressing vacuum decay as found in [34], a point to which we return later in the context of anti-confining quenches at the end of Section 5.
@@ 4 Escaping fronts in confining local quenches
Naively, the escaping fronts observed in the local quenches looked as if a meson pair had traced them out. However, a simple energy estimate shows that this cannot be correct. The quench is generated by flipping a local spin, which does not inject enough energy into the system. For a crude estimate, note that the main contribution comes from the nearest neighbour interaction in (2.1), which can contribute at most $4J$ for a single spin flip. The contribution by the longitudinal and transverse fields cannot be larger than $2J(h_{l}+h_{t})$, so altogether, the injected energy is smaller than $(4+2h_{l}+2h_{t})J$. For all the quenches considered in Subsection 2.2, this is never enough to cover the energy of a meson pair. For example, taking $h_{t}=0.25$ and $h_{l}=0.1$, the injected energy is estimated from above by $5.2J$. At the same time, the 2-fermion approximation described in Section 3 yields $m_{1}=3.668J$ for the energy of the lightest meson with momentum zero, i.e. the meson gap. Therefore, the contribution of states with more than one meson is expected to be negligible.
However, the above considerations suggest that the explanation of the escaping fronts must be sought in terms of single-meson states. To reveal the nature of the states that contribute, we analyse the overlaps of the initial state with the post-quench meson states.
@@ 4.3 Meson overlaps
  • The state (4.2), constructed neglecting the longitudinal field, can be considered an excellent approximation of the actual initial state. In particular, the actual initial state also lies predominantly in the two-fermion subspace. As a result, the expansion is dominated by single-meson states.
@@
These features allow a semi-analytic calculation where the initial state is approximated by the analytic expression derived in Subsection 4.1, while the meson states are computed in the two-fermion approximation by solving the Schrödinger equation (3.2). The latter, while non-analytic, can be performed to any desired accuracy; therefore, the final overlaps can be considered free of numerical errors. Additionally, the method allows for the overlaps to be computed directly for a chain of any length, making the results applicable in the thermodynamic limit of an infinite system. In practice, we computed the overlaps for a chain of length $L=100$, which is effectively nothing else but a discretisation of the momentum integral corresponding to the infinite volume limit.
@@
which has a small global component. An example of such a quench is presented in Fig. C.1 a) of Appendix C, for the parameters $h_{t}=0.25$ and $h_{l}=0.2$. To the naked eye, there is no obvious difference between the time evolutions of the longitudinal magnetisation shown in Fig. C.1 a) and in the purely local quench shown in the second panel of Fig. 2.1. Indeed, the deviation turns out to be suppressed by around $3$ orders of magnitude compared to the signal from the localised component of the purely local quench.
@@ 5 Escaping fronts in anti-confining local quenches
where the relative sign between the NS and R components corresponds to the state built by a spin flip performed in the false vacuum as discussed at the end of Subsection 4.1. We note that the bubble wave functions in momentum space contain a similar relative sign between the NS and R components, again reflecting that these excitations are built upon the false vacuum, and these two signs eventually cancel when computing the overlaps
@@
To sum up, for the anti-confining quenches, dynamical confinement is replaced by Wannier-Stark localisation, but again, it is the Schrödinger cats which escape it, with the only difference is that these are bubble cats instead of mesonic ones.
@@ 6 Conclusions and outlook
We also considered anti-confining quenches initiated by a single spin flip in the false vacuum. In global quenches, the light cone spreading of correlations is again suppressed; however, this time by the mechanism of Wannier-Stark localisation [28] based on Bloch oscillations, which originate from the periodicity of the lattice dispersion relation in momentum space, and also result in the suppression of the decay of the false vacuum [34]. In the local quenches considered here, we once again observed a significant propagating component of magnetisation besides the localised one. This time, the excitations responsible for the fronts are bubbles of the true vacuum nucleating from the false vacuum, again in superpositions of left and right-moving bubbles (dominated by the one with the smallest spatial extension) with equal weights. Interestingly, in infinite volume, the spectrum of these bubbles is bounded from above while unbounded from below. In contrast, in a finite volume, their spectrum seamlessly blends into that of the mesons since both large bubbles and large mesons correspond to collisionless kink pairs separated by distances much larger than the Wannier-Stark localisation length.
@@
It would be interesting to perform a quantum simulation of these phenomena like those already attempted for global quenches on a quantum computer or in a suitably designed ultra-cold atomic experiment. Another interesting direction is to consider these phenomena in the framework of the quantum $3$-state Potts spin chain in its confining phase, where the spectrum contains baryonic excitations besides the mesonic ones [55, 56, 57, 58], strengthening the analogy with the strong interactions and enlarging the scope of potential physical phenomena.
@@ Acknowledgments
We are grateful to M.A. Werner for sharing his extensive knowledge of MPS methods and helping us develop our simulations. We also thank M. Kormos and M. Lencsés for discussions, and the latter especially for suggesting the title of our work. This work was supported by the National Research, Development and Innovation Office (NKFIH) through the OTKA Grants SNN 139581 and K 138606. AK was also partially supported by the Doctoral Excellence Fellowship Programme (DCEP), funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Office. GT was partially supported by the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004).
@@ Appendix B Numerical results for meson and bubble wave functions
modified by the (inverse of the) phase factors specified in (4.19). This results in an excellent match, as illustrated in Fig. B.1 for mesons and Fig. B.2 for bubbles, even on a chain of moderate length.
@@ Appendix C Time evolution after other local quenches
The suppression of escaping fronts in the domain wall quench is in marked contrast to the quenches initialised by local spin flips considered in Section 2. This can be understood on the confining side of the quench by reconsidering the energy balance argument presented at the beginning of Section 4. The energy locally fed into the system by the presence of the domain wall is roughly smaller by a factor of $1/2$ due to the presence of a single boundary, instead of the two which occur at a spin flip and is therefore insufficient for the creation of even the lowest lying meson state. However, this argument is insufficient as an explanation because, on the anti-confining side, the bubble spectrum is unbounded from below, seemingly allowing for the unlimited creation of bubbles. Nevertheless, note that bubbles with small spatial extensions (i.e., species index) have positive energy and only supercritical bubbles with sizes larger than $n_{*}$ in Eqn. (3.20) can turn the energy balance in a favourable direction. As discussed in Section 5, the size of these bubbles is so large that they have extremely small overlaps with the initial state. Additionally, such supercritical bubbles are collisionless and therefore stationary, as discussed in Section 3.

### 24 Jan 2024 v2

@@ Abstract
We consider local quenches from initial states generated by a single spin-flip in either the true or the false vacuum state of the confining quantum Ising spin chain in the ferromagnetic regime. Contrary to global quenches, where the light-cone behaviour is strongly suppressed, we find a significant light-cone signal propagating with a nonzero velocity besides the expected localised oscillating component. Combining an analytic representation of the initial state with a numerical description of the relevant excitations using the two-fermion approximation, we can construct the spectrum of post-quench excitations and their overlaps with the initial state, identifying the underlying mechanism. For confining quenches built upon the true vacuum, the propagating signal consists of Schrödinger cats of left and right-moving mesons escaping confinement. In contrast, for anti-confining quenches built upon the false vacuum, it is composed of Schrödinger cats of left and right-moving bubbles which escape Wannier-Stark localisation.
@@ 1 Introduction
It was shown in [3] that the non-equilibrium dynamics of the quantum Ising spin chain is radically altered by confinement. In this context, non-equilibrium time evolution was initiated by a quantum quench, i.e., a sudden change in the Hamiltonian [4, 5]. This protocol is of paradigmatic relevance, and today, it is routinely engineered in experiments on closed quantum systems [6, 7, 8, 9, 10, 11, 12, 13]. In a large class of such systems, including various spin chains with short-range interactions or interacting bosonic or fermionic systems, quasi-particles have an upper bound on their velocity [14], leading typically to a distinctive light-cone pattern in time-dependent correlation functions [15, 4, 5, 16, 17, 18, 19, 20]. However, as shown in [3], the confining forces can suppress the light-cone spreading of correlations. Dynamical confinement and its effects have recently been confirmed in numerous systems  [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@ 2.1 Confinement in the Ising spin chain
$\chi=2Jh_{l}m\,$ (2.4)
between kinks, which confines them into meson excitations. In the scaling limit, the disappearance of kinks from the spectrum can also be explained in field-theoretic terms based upon the mutual non-locality of the order parameter and the interpolating field of the kinks [51].
Choosing for definiteness $m>0$, the true ground state of the system is a perturbation of the one with positive magnetisation $\ket{+}$, while $\ket{-}$ becomes a metastable state, a.k.a. the false vacuum. In the following, we stay within the scenario of strong confinement, which corresponds to a transverse field $h_{t}$ far below its critical value $1$ and consider values of the longitudinal field $h_{l}$ of order $0.1$.
@@
where $\hat{k}_{1,2}$ and $\hat{x}_{1,2}$ are the canonical momentum and coordinate operators for the kinks. The two-kink state dominance of the meson states was explicitly demonstrated in the field-theoretic limit [53], and it is related to the suppression of kink number changing processes [54].
The spectrum of this Hamiltonian can be computed either by a numerical solution of the Schrödinger equation or using a simple semi-classical approximation [52]. We avoid using values that are too large for $h_{l}$ to ensure that even the most spatially localised meson excitations have support over several lattice sites, which makes it possible to neglect the short-range deviations from the linear potential. These were computed in [52] and result from the kinks not being exactly point-like: the interpolation between the two vacua takes place over a crossover region of finite length.
@@
We note that when the polarisation of the initial state is opposite to the longitudinal field present after the quench, the time evolution is again localised in the global case; however, in this case, the mechanism behind this effect is Wannier-Stark localisation due to Bloch oscillations [26]. This corresponds to starting the dynamics from a false vacuum state, whose decay is prevented by Bloch oscillations [38]. Contrary to confinement, which also occurs in the continuum limit [53], Wannier-Stark localisation is a lattice effect resulting from the dispersion relation’s periodicity over the Brillouin zone.
@@ 2.2 Phenomenology of local quenches
The first quench protocol we consider is the confining quench starting from the initial state
@@
As a first step, the system’s ground state was found by applying the TEBD method with imaginary time evolution and was validated by comparing it to the exact diagonalisation results. We then applied the spin flip and simulated the real-time evolution of the system. For the imaginary time evolution, we chose the time step $dt=0.05$, the total number of steps $N=4000$ and the maximum bond dimension $\chi_{\text{max}}=50$. For the real-time evolution, the parameters were set to $dt=0.01$ and $\chi_{\text{max}}=300$. The Schmidt values were truncated at $10^{-8}$ for both imaginary and real-time evolutions. We verified that all results were robust under reasonable variations of these parameters.
@@
Note that while the localised component qualitatively matches the predicted envelope, there is a component forming a prominent light cone which lies outside the red curves. In contrast to the results for global quenches [3], the escaping fronts in local quenches are of the same order of magnitude as the localised component. Similar dynamics was observed in [31]; however, the transverse field applied there was much larger ($h_{t}=0.5$), much closer to the critical point. They also considered a different observable: the probability dynamics of kinks instead of the magnetisation which is considered here. It is striking to see such a prominent light-cone signal in the strongly confining regime where the light-cone in global quenches was observed to be very strongly suppressed [3]. Our primary interest is to identify the mechanism underlying their presence.
@@
where |$\ket{+}_{h_{t},h_{l}=0}$ is the positively polarised ground state of the pure transverse Hamiltonian. In this case, we take the post-quench longitudinal field $h_{l}$ to be negative, and so the initial state corresponds to a metastable (false vacuum) state for the Hamiltonian (2.1). The general settings of the TEBD simulations for these quenches were the same as in the confining case. Note that such quenches are not purely local, i.e. they necessarily involve a global component; however, it turns out that this component is so small that it can be neglected in our subsequent considerations.
@@ 3.1 Confinement: mesons
where $\omega(\hat{k};K)=\varepsilon(K/2+\hat{k})+\varepsilon(K/2-\hat{k})$. From now on, we will omit the hats from the operators. Considering an infinite chain with position $x$ taking values in $\mathbb{Z}$, the Schrödinger equation can be written as
@@
The $n$th level of the above Schrödinger problem corresponds to a meson one-particle state $|M_{n}(K)\rangle$ with momentum $K$ and species number $n$, with $E_{n}(K)$ giving its energy relative to the ground state in the two-fermion approximation, which we order to increase with $n$. For any value of $K$, the meson spectrum is bounded from below and unbounded from above.
@@
$d_{\text{WS}}=\frac{\epsilon(\pi)-\epsilon(0)}{\chi}=\frac{2h_{t}}{\chi}\,,$ (3.8)
@@
up to a phase factor. For the collisionless case, imposing periodicity of the wave function in the Brillouin zone as ${\psi}(k=\pi)={\psi}(k=-\pi)$ leads to a Bohr-Sommerfeld quantisation condition of the form
@@
This can be checked to match numerically the wave function obtained by the explicit solution of (3.2) for large $n$. For the case $h_{t}=0.25$ and $h_{l}=0.1$, the agreement is already excellent for $n=6$ and rapidly improves for larger $n$.
@@ 3.2 Anti-confinement: bubbles
which grows monotonically with $h_{t}$, and is bounded from below by $2/|h_{l}|$. For the transverse and longitudinal fields considered here, the critical bubble size is always more than one order of magnitude larger than the lattice spacing, which plays an important role in suppressing vacuum decay as found in [38], a point to which we return later in the context of anti-confining quenches at the end of Section 5.
@@ 4 Escaping fronts in confining local quenches
Naively, the escaping fronts observed in the local quenches looked like as if a meson pair had traced them out. However, a simple energy estimate shows that this cannot be correct. The quench is generated by flipping a local spin, which does not inject enough energy into the system. For a crude estimate, note that the main contribution comes from the nearest neighbour interaction in (2.1), which can contribute at most $4J$ for a single spin-flip. The contribution by the longitudinal and transverse fields cannot be larger than $2J(h_{l}+h_{t})$, so altogether, the injected energy is smaller than $(4+2h_{l}+2h_{t})J$. For all the quenches considered in Subsection 2.2, this is never enough to cover the energy of a meson pair. For example, taking $h_{t}=0.25$ and $h_{l}=0.1$, the injected energy is estimated from above by $5.2J$. At the same time, the 2-fermion approximation described in Section 3 yields $m_{1}=3.668J$ for the energy of the lightest meson with momentum zero, i.e. the meson gap. Therefore, the contribution of states with more than one meson is expected to be negligible.
The above considerations suggest that the explanation of the escaping fronts must be sought in terms of single-meson states. To reveal the nature of the states that contribute, we analyse the overlaps of the initial state with the post-quench meson states.
@@ 4.3 Meson overlaps
  • The state (4.2), constructed neglecting the longitudinal field, can be considered an excellent approximation of the actual initial state. In particular, the exact initial state also lies predominantly in the two-fermion subspace. As a result, the expansion is dominated by single-meson states.
@@
These features allow a semi-analytic calculation where the initial state is approximated by the analytic expression derived in Subsection 4.1, while the meson states are computed in the two-fermion approximation by solving the Schrödinger equation (3.2). The latter, while non-analytic, can be performed to any desired accuracy; therefore, the final overlaps can be considered free of numerical errors. Additionally, the method allows for the overlaps to be computed directly for a chain of any length, making the results applicable in the thermodynamic limit of an infinite system. In practice, we computed the overlaps for a chain of length $L=100$, which effectively discretises the momentum integral corresponding to the infinite volume limit.
@@
which has a small global component. An example of such a quench is presented in Fig. C.1 a) of Appendix C, for the parameters $h_{t}=0.25$ and $h_{l}=0.2$. To the naked eye, there is no obvious difference between the time evolutions of the longitudinal magnetisation shown in Fig. C.1 a) and of the purely local quench shown in the second panel of Fig. 2.1. Indeed, the deviation turns out to be suppressed by around $3$ orders of magnitude compared to the signal from the localised component of the purely local quench.
@@ 5 Escaping fronts in anti-confining local quenches
where the relative sign between the NS and R components corresponds to the state built by a spin-flip performed in the false vacuum as discussed at the end of Subsection 4.1. We note that the bubble wave functions in momentum space contain a similar relative sign between the NS and R components, again reflecting that these excitations are built upon the false vacuum, and these two signs eventually cancel when computing the overlaps
@@
To sum up, for the anti-confining quenches, dynamical confinement is replaced by Wannier-Stark localisation, but again, it is the Schrödinger cats which escape it, with the only difference that these are bubble cats instead of mesonic ones.
@@ 6 Conclusions and outlook
We also considered anti-confining quenches initiated by a single spin-flip in the false vacuum. In global quenches, the light cone spreading of correlations is again suppressed; however, this time by the mechanism of Wannier-Stark localisation [26] based on Bloch oscillations, which originate from the periodicity of the lattice dispersion relation in momentum space, and also result in the suppression of the decay of the false vacuum [38]. In the local quenches considered here, we once again observed a significant propagating component of magnetisation besides the localised one. This time, the excitations responsible for the fronts are bubbles of the true vacuum nucleating from the false vacuum, again in superpositions of left and right-moving bubbles (dominated by the one with the smallest spatial extension) with equal weights. Interestingly, in infinite volume, the spectrum of these bubbles is bounded from above while unbounded from below. In contrast, in a finite volume, their spectrum seamlessly blends into that of the mesons since both large bubbles and large mesons correspond to collisionless kink pairs separated by distances much larger than the Wannier-Stark localisation length.
@@
It is also interesting to perform a quantum simulation of these phenomena on a quantum computer or in a suitably designed ultra-cold atomic experiment. While confining local quenches have previously been considered on a quantum computer [31], we are unaware of similar results for anti-confining quenches. Another interesting direction is to consider these phenomena in the framework of the quantum $3$-state Potts spin chain in its confining phase, where the spectrum contains baryonic excitations besides the mesonic ones [60, 61, 62, 63], strengthening the analogy with the strong interactions and enlarging the scope of potential physical phenomena.
@@ Acknowledgments
We are grateful to M.A. Werner for sharing his extensive knowledge of MPS methods and helping us develop our simulations. We also thank M. Kormos and M. Lencsés for discussions, and the latter especially for suggesting the title of our work. We are grateful to A. Bastianello for useful comments on the initial version of the manuscript. This work was supported by the National Research, Development and Innovation Office (NKFIH) through the OTKA Grants SNN 139581 and K 138606. AK was also partially supported by the Doctoral Excellence Fellowship Programme (DCEP), funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Office. GT was partially supported by the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004).
@@ Appendix B Numerical results for meson and bubble wave functions
modified by the (inverse of the) phase factors specified in (4.19). This results in an excellent match, as illustrated in Fig. B.1 for mesons and in Fig. B.2 for bubbles, even on a chain of moderate length.
@@ Appendix C Time evolution after other local quenches
The suppression of escaping fronts in the domain wall quench is in marked contrast to the quenches initialised by local spin-flips considered in Section 2. This can be understood on the confining side of the quench by reconsidering the energy balance argument presented at the beginning of Section 4. The energy locally fed into the system by the presence of the domain wall is roughly smaller by a factor of $1/2$ due to the presence of a single boundary, instead of the two which occur at a spin-flip and is therefore insufficient for the creation of even the lowest lying meson state. However, this argument is insufficient as an explanation because, on the anti-confining side, the bubble spectrum is unbounded from below, seemingly allowing for the unlimited creation of bubbles. Nevertheless, note that bubbles with small spatial extensions (i.e., species index) have positive energy and only supercritical bubbles with sizes larger than $n_{*}$ in Eqn. (3.20) can turn the energy balance in a favourable direction. As discussed in Section 5, the size of these bubbles is so large that they have extremely small overlaps with the initial state. Additionally, such supercritical bubbles are collisionless and therefore stationary, as discussed in Section 3.